OpenStudy (anonymous):

Using the following equation, find the center and radius of the circle. You must show all work and calculations to receive credit. x2 + 4x + y2 - 6y = -4 (8 points)

OpenStudy (anonymous):

Can someone please help me

OpenStudy (solomonzelman):

\(\LARGE\color{black}{ \bf x^2 + 4x + y^2 - 6y = -4}\) \(\LARGE\color{black}{ \bf x^2 + 4x\color{red} { +4 } + y^2 - 6y\color{blue} { +9 } = -4\color{red} { +4}\color{blue} { +9 } }\)

OpenStudy (solomonzelman):

Can you finish the problem off, knowing that the equation of the circle looks like, \(\LARGE\color{black}{ \bf (x-h)^2+(y-k)^2=r^2}\)

OpenStudy (solomonzelman):

Where (h,k) is the center, and r is the radius.

OpenStudy (solomonzelman):

I am giving one more hint. \(\LARGE\color{red}{ \bf x^2+4x+4}\) and \(\LARGE\color{blue}{ \bf y^2-6y+9}\) are both perfect squares.

OpenStudy (acxbox22):

*trinomials

OpenStudy (anonymous):

I think i have the center correct but im a little confused on finding the radius @SolomonZelman @acxbox22

OpenStudy (anonymous):

Oh never mind i got it now